# A Scalable Pipeline for Enabling Non-Verbal Speech Generation and Understanding

### v1

@@ 8 Conclusion
In this work, we present NonVerbalSpeech-38K, a large, diverse dataset for advancing non-verbal TTS and end-to-end speech captioning. By collecting naturally expressive audio from diverse media and developing an automated annotation pipeline, it enhances current speech synthesis and understanding systems in non-verbal generation and understanding, respectively—with its effectiveness validated by experiments using state-of-the-art models.
Appendix A Experimental Setup
@@
Dataset Laughing Coughing Breath Total
CapSpeech (Train) 19,571 15,174 2,110 36,855
NonVerbalSpeech-38K Subset (Train) 6,912 5,940 2,162 15,014
NonVerbalTTS Subset (Eval) 367 137 250 754
Table 7: Counts of non-verbal tags in training and evaluation sets. CapSpeech and a subset of NonVerbalSpeech-38K are used for training, and the NonVerbalTTS subset is used for evaluation.
@@ A.1 Experimental Setup and Results for Frame-level Detection Model
We set $N_{q}$ to 10, corresponding to the 10 non-verbal phenomena described in Table 2, and the number of layers $L$ for the pre-trained wav2vec 2.0 model (Baevski et al. 2020) is 25. Both the encoder and decoder are configured with 6 layers, where the decoder excludes the self-attention module. The $Q$ embeddings are randomly initialized. We adopt LoRA (Hu et al. 2022) to fine-tune all linear layers within the self-attention modules of the wav2vec 2.0 backbone. The model is trained for 300 epochs with a batch size of 64 per GPU, using 8 NVIDIA H100 GPUs. The learning rate is set to $1\times 10^{-4}$ and scheduled with a warm-up phase followed by cosine annealing decay.
@@
Table 8: Example alignment between reference (REF) and model hypothesis (HYP) with non-verbal tag [coughing]. Insertions (e.g., repeated what, after) and substitutions (e.g., he vs. they) are observed.
A.2 Experimental Setup for Non-Verbal Speech Generation
Implementation Details
In the non-verbal speech generation task, we use F5-TTS (Chen et al. 2024) as the base model. It is a fully non-autoregressive speech synthesis model that adopts a flow-matching method combined with a transformer architecture. We use the pretrained F5-TTS model, which was trained on the Chinese and English subsets of the Emilia dataset (He et al. 2024), and further fine-tune it on our target data. We use BigVGAN-v2 (Lee et al. 2022) as the vocoder. BigVGAN-v2 is trained on datasets that include diverse types of audio, such as speech in multiple languages, environmental sounds, and musical instruments. This diversity is important for non-verbal speech generation, as it helps the vocoder better handle a wide range of acoustic patterns.
To support non-verbal control, we expand the vocabulary of F5-TTS by adding non-verbal tags such as [laughing], [coughing]. Inspired by CapSpeech (Wang et al. 2025), we initialize the embeddings of the new tokens using representations from the CLAP (Elizalde et al. 2023) text encoder, which is trained on large-scale audio-text alignment data, including non-verbal audio-text pairs.
We fine-tune F5-TTS on three datasets: CapSpeech (Wang et al. 2025), NonVerbalTTS (Borisov, Spirin, and Diatlova 2025), and our proposed NonVerbalSpeech-38K dataset. To retain the zero-shot synthesis ability of the pretrained model while allowing it to learn non-verbal speech generation, we carefully adjust the total number of training steps according to each dataset. Based on empirical observations, we train both NonVerbalTTS and our proposed NonVerbalSpeech-38K dataset for 400 epochs, while CapSpeech is only trained for 50 epochs. We find that training CapSpeech for 400 epochs significantly degrades intelligibility, making the synthesized speech hard to understand. All models are optimized using the AdamW optimizer with a base learning rate of $1\mathrm{e}{-5}$, a 10% warm-up phase, and a cosine annealing learning rate schedule. For the newly added non-verbal tokens, a higher learning rate of $1\mathrm{e}{-4}$ is applied to encourage faster adaptation. Fine-tuning is performed on 8 NVIDIA H100 GPUs, with a batch size of 38,400 frames per GPU.
Metrics QMOS: Quality Evaluation IMOS: Instruction Following Evaluation
Guidance Please rate each audio sample based on naturalness (whether it sounds like real human speech), sound quality (whether it is clean and undistorted), and intelligibility (how easily the content can be understood). Note: Segments containing only silence should be excluded from the evaluation of naturalness and intelligibility. Please rate each audio sample based on how well it follows the given text instructions. For example, given the text: [sigh] Today’s work is not even halfway done yet.. The TTS system is expected not only to synthesize the text correctly but also to follow the [sigh] tag by producing a sigh sound at the corresponding position. All tags used in the evaluation and their meanings are as follows: [laughing]: laughing sound, [coughing]: coughing sound, [sigh]: sighing sound, [sniff]: sniffing sound, [throatclearing]: throat clearing sound, [breath]: breathing sound.
Score Criteria Criteria
5 Excellent naturalness, pure sound quality, and perfectly intelligible content. Fully follows the text and accurately renders the non-verbal tags, the speech sounds entirely human-like.
4 Good naturalness, mostly clean sound, and clear content. Fully follows the text, and non-verbal tags are rendered naturally.
3 Moderate naturalness, minor sound artifacts (e.g., slight noise), and mostly intelligible content. Mostly follows the text; non-verbal tag rendering is generally accurate but lacks naturalness.
2 Poor naturalness, noticeable noise or distortion, and hard-to-understand content. Slight omission of text, missing or incorrectly rendered tags (e.g., misplaced tag or wrong sound).
1 Very unnatural (e.g., mechanical), heavily distorted, and nearly unintelligible. Significant text loss or synthesis failure (e.g., unintelligible or missing content).
Table 9: Subjective evaluation criteria for QMOS and IMOS.
Evaluation
To evaluate system performance, in addition to standard metrics such as DNSMOS, SSIM (speaker similarity), and WER/CER (Word Error Rate or Character Error Rate) for assessing speech quality, speaker similarity, and intelligibility, respectively, we also introduce the CLAP Score to evaluate whether the synthesized speech contains the intended non-verbal phenomena. Specifically, we use the sentence “This is a sound of non-verbal-tag.” as the text input to the CLAP (Elizalde et al. 2023) text tower, and feed the full synthesized waveform into the CLAP audio tower. For subjective evaluation, we adopt two metrics. The first is QMOS, which assesses the overall quality of the synthesized speech, focusing on intelligibility, audio quality, and naturalness. The second is IMOS, which evaluates whether the audio follows the input text, including non-verbal tags. For example, given the input “[sigh] The work for today is not even halfway done yet.”, the TTS system is expected not only to correctly synthesize the verbal content but also to emit a sigh sound at the appropriate position as indicated by the [sigh] tag. We collected a total of 20 subjective questionnaires. Detailed scoring criteria are provided in the table 9.
@@ A.2 Experimental Setup for Non-Verbal Speech Generation
Listing 1 Insert Non-Verbal Tag into Transcription
1# Inputs:
2# n: [n_s, n_e], non-verbal segment
3# l: non-verbal label
4# W: word list [(s_i, e_i, w_i)] for i = 1 to N
5s_tag = f"[{l}]<B>"
6e_tag = f"</B>"
7# Find word indices overlapping with the non-verbal segment
8i_s = max(i for i, (s, _, _) in enumerate(W) if s <= n_s)
9i_e = min(i for i, (_, e, _) in enumerate(W) if e >= n_e)
10if i_e - i_s > 1:
11 # Spans multiple words
12 return W[:i_s + 1] + [s_tag] + W[i_s+1:i_e] + [e_tag] + W[i_e:]
13else:
14 # Single word
15 return W[:i_e+1] + [f"[{l}]"] + W[i_e+1:]
A.3 Experimental Setup for End-to-End Non-Verbal Speech Understaning
Implementation Details
In this task, we use Qwen2-Audio (Chu et al. 2024) as our base model. The fine-tuning is performed using LoRA, applied to the linear layers within the self-attention modules of the language model. During training, only the LoRA parameters and the newly introduced tokens are updated, while all other model parameters are frozen. For a fair comparison, we use only the subset of NonVerbalSpeech-38K that contains the three non-verbal tags supported by CapSpeech (i.e., [laughing], [breath], and [coughing]), the details are shown in Table 7. Both datasets are trained for 5 epochs. All models are trained using a learning rate of 7.5e-5, batch size 12 per GPU, with 10% warmup and cosine learning rate decay, using 8 NVIDIA H100 GPUs.
Evaluation
The subset of NonVerbalTTS for evaluation is shown in Table 7. To simplify the evaluation, we include only samples that contain a single non-verbal tag, resulting in 367 [laughing], 137 [coughing], and 1,625 [breath] samples. To ensure a balanced tag distribution, 250 samples are randomly selected from the 1,625 available [breath] samples. The resulting evaluation set accounts for approximately 5% of the NonVerbalSpeech-38K subset used for training.
For the metrics, in addition to the standard word error rate (WER), we introduce three additional metrics focusing on the correctness of predicted tags in terms of both type and position:
  • F1-Score: Measures the correctness of predicted tags compared to the ground truth.
    $\text{TPD}=\frac{1}{n}\sum_{i=1}^{n}|p_{i}-g_{i}|.$ (4)
Appendix B NonVerbalSpeech-38K Pipeline
B.1 Details on Incorporating Non-Verbal Tags into Speech Content
To support non-verbal speech generation and understanding, it is necessary to integrate the textual content with non-verbal tags. The related algorithm is shown in the pseudocode 1. It is worth noting that state-of-the-art ASR models—including Whisper-large-v3 (Radford et al. 2023) and Paraformer-zh (Gao et al. 2022)—offer word-level timestamps. In practice, however, we have observed that these timestamps exhibit varying degrees of misalignment, with Paraformer-zh showing particularly noticeable discrepancies. We try using forced alignment tools, such as Montreal Forced Aligner (McAuliffe et al. 2017), to improve alignment. But due to the high WER in the transcriptions, no significant improvement is achieved, in some cases, the results are even worse.
@@ Appendix C NonVerbalSpeech-38K Dataset
Using the proposed NonVerbalSpeech-38K pipeline, we construct the NonVerbalSpeech-38K dataset from a large collection of speech data crawled from diverse online sources. This diversity ensures coverage of a wide range of natural speaking styles and, importantly, includes rich and varied non-verbal phonemes. The resulting durations mostly fall within the range of 3 to 30 seconds (Figure 5), primarily because the Emilia (He et al. 2024) pipeline segments speech clips into chunks of 3–30 seconds. This segmentation operation also explains the peak near 30 seconds. Some segments slightly exceeding 30 seconds arise during Step 3 of the NonVerbalSpeech-38K pipeline (Algorithm 1), where non-verbal segments that do not fall within any speech region but are within a 1-second distance threshold are merged with the nearest speech region to maintain contextual continuity.

### v2

@@ 7 Conclusion
In this work, we design a fully automated NV annotation framework featuring two novel components—a unified detection model and the TSA alignment module—avoiding reliance on costly manual labeling and the unnatural outputs caused by rule-based synthesis in prior work. Using this framework, we directly annotate NV phenomena from in-the-wild speech, creating and publicly releasing the NVS dataset, which contains diverse and natural NV events with precise textual labels. Experimental results demonstrate that this dataset effectively enabling existing models to perform NV generation and recognition, particularly exhibiting superior NV controllability.
Appendix A Experimental Setup and Results
@@ A.1 Experimental Setup and Results for Frame-level Detection Model
We set $N_{q}$ to 10, corresponding to the 10 non-verbal phenomena, and the number of layers $L$ for the pre-trained wav2vec-bert 2.0 model (Barrault et al. 2023) is 25. Both the encoder and decoder are configured with 6 layers, where the decoder excludes the self-attention module. The $Q$ embeddings are randomly initialized. We adopt LoRA (Hu et al. 2022) to fine-tune all linear layers within the self-attention modules of the wav2vec-bert 2.0 model. The model is trained for 300 epochs with a batch size of 64 per GPU, using 8 NVIDIA H100 GPUs. The learning rate is set to $1\times 10^{-4}$ and scheduled with a 5% warm-up phase followed by cosine annealing decay.
@@
A.2 Experimental Setup and Results for Non-Verbal Speech Generation
Implementation Details
In the non-verbal speech generation task, we use F5-TTS (Chen et al. 2024) as the base model. It is a fully non-autoregressive speech synthesis model that adopts a flow-matching method combined with a transformer architecture. We use the pretrained F5-TTS model, which was trained on the Chinese and English subsets of the Emilia dataset (He et al. 2024), and further fine-tune it on the non-verbal dataset. We use BigVGAN-v2 (Lee et al. 2022) as the vocoder. BigVGAN-v2 is trained on datasets that include diverse types of audio, such as speech in multiple languages, environmental sounds, and musical instruments. This diversity is important for non-verbal speech generation, as it helps the vocoder better handle a wide range of acoustic patterns. To support non-verbal control, we expand the vocabulary of F5-TTS by adding non-verbal tags such as [laughing], [coughing]. Inspired by CapSpeech (Wang et al. 2025), we initialize the embeddings of the new tokens using representations from the CLAP (Wu et al. 2023) text encoder, which is trained on large-scale audio-text alignment data, including non-verbal audio-text pairs.
To retain the basic zero-shot text-to-speech ability of the pretrained model while allowing it to learn non-verbal speech generation, we carefully adjust the total number of training steps based on the scale of each dataset. Based on empirical observations, for all datasets—including ours—but excluding NVSpeech (Liao et al. 2025), SynParaSpeech (Bai et al. 2025), and CapSpeech (Wang et al. 2025), we train the models for 400 epochs to avoid overfitting and maintain fundamental TTS capability. For CapSpeech, we train the model for 50 epochs to achieve the best performance, as extending the training to 400 epochs significantly degrades intelligibility and results in noticeably unnatural speech. NVSpeech and SynParaSpeech are trained for approximately 60k steps, matching the number of training steps used for our dataset, which ensures a fair comparison while preventing overfitting.
All models are optimized using the AdamW optimizer with a base learning rate of $1\mathrm{e}{-5}$, a 10% warm-up phase, and a cosine annealing learning rate schedule. For the newly added non-verbal tokens, a higher learning rate of $1\mathrm{e}{-4}$ is applied to encourage faster adaptation. Fine-tuning is performed on 8 NVIDIA H100 GPUs, with a batch size of 38,400 frames per GPU.
Evaluation
To evaluate system performance, in addition to standard metrics such as DNSMOS [^3] https://github.com/microsoft/DNS-Challenge, SSIM (speaker similarity) [^4] https://github.com/BytedanceSpeech/seed-tts-eval, and WER/CER (Word Error Rate or Character Error Rate) [^5] https://github.com/BytedanceSpeech/seed-tts-eval for assessing speech quality, speaker similarity, and intelligibility, respectively, we also introduce the CLAP Score (Wu et al. 2023) to evaluate whether the synthesized speech contains the intended non-verbal phenomena. Specifically, we use the template sentence “This is a sound of specific non-verbal sound.” (e.g., for laughing, “This is a sound of laughing.”) as the text input to the CLAP (Wu et al. 2023) text tower, and feed the entire synthesized waveform into the CLAP audio tower. For subjective evaluation, we adopt two metrics. The first is QMOS, which assesses the overall quality of the synthesized speech, focusing on intelligibility, audio quality, and naturalness. The second is IMOS, which evaluates whether the audio follows the input text, including non-verbal tags. For example, given the input “[sigh] The work for today is not even halfway done yet.”, the TTS system is expected not only to correctly synthesize the verbal content but also to emit a sigh sound at the appropriate position as indicated by the [sigh] tag. We collected a total of 20 subjective questionnaires. Detailed scoring criteria are provided in the table 6.
Experimental Results
Table 9 presents IMOS comparisons across NV tags. We observe that NVS (TSA) ranks below existing models in only a few cases and achieves consistent leadership in Chinese, demonstrating comprehensive and stable NV controllability.
A.3 Experimental Setup for Non-Verbal Speech Understaning
Implementation Details
In this task, we use Qwen2-Audio and Whisper-Large-V3 as base models, extending their vocabularies to support non-verbal tags (e.g., [laughing], [breath]). For Qwen2-audio, the fine-tuning is performed using LoRA, applied to the linear layers within the self-attention modules of the language model. During training, only the LoRA parameters and the newly introduced tokens are updated, while all other model parameters are frozen. All models are trained using a learning rate of 7.5e-5, batch size 12 per GPU, with 10% warmup and cosine learning rate decay, using 8 NVIDIA H100 GPUs. For Whisper-Large-V3, we full-parameters fine-tuneing the whisper decoder while keep the encoder frozen. All models are trained using a learning rate of 7.5e-5, batch size 24 per GPU, with 10% warmup and cosine learning rate decay, using 8 NVIDIA H100 GPUs.
In the English setting, for a fair comparison, we use only the subset of NonVerbalSpeech-38K containing the three non-verbal tags supported by CapSpeech (i.e., [laughing], [breath], and [coughing]) for training. Details are shown in Table 7. Both datasets (including the two variants of our NonVerbalSpeech-38K) are trained for 5 epochs. In the Chinese setting, we use the overlapping tags in NVSpeech and our dataset (i.e., [laughing], [coughing], [sigh]) for training. Details are shown in Table 8. Both datasets (including the two variants of our NonVerbalSpeech-38K) are trained for 5 epochs. Note that both CapSpeech and NVSpeech have more total samples than our NVS dataset. Even under this setting, our dataset outperforms CapSpeech and achieves performance comparable to NVSpeech, demonstrating its efficiency and effectiveness.
Evaluation
In the English setting, we use a subset of NonVerbalTTS for evaluation, as shown in Table 7. In the Chinese setting, a subset of SMIIP-NV and a subset of MNV-17 are combined and used for evaluation, as shown in Table 8. To simplify the evaluation, we include only samples that contain a single non-verbal tag.
For the evaluation metrics, in addition to standard WER/CER (excluding non-verbal tags) and NV-aware WER/CER (NVWER/NVCER, including them), we introduce two tag-level metrics to further evaluate position accuracy: Tag Position Distance (TPD) and its normalized variant, Normalized Tag Distance (NTD).
    $\text{TPD}=\frac{1}{n}\sum_{i=1}^{n}|p_{i}-g_{i}|.$ (5)
Table 9: IMOS Differences (Comparison Systems - F5-TTS Fine-Tuned with NVS (TSA)), Broken Down by Non-Verbal Tag.
Model [laughing] [throat] [sigh] [coughing] [sniff] [breath]
English (EN)
CosyVoice2 1.40-1.40 - 1.08-1.08 1.60-1.60 - 1.15-1.15
Dia $\mathbf{0.22}$ $\mathbf{0.53}$ 0.95-0.95 1.37-1.37 0.40-0.40 -
F5-TTS - - - - - -
+ CapSpeech 1.81-1.81 - - 2.03-2.03 - 1.38-1.38
+ NVTTS 0.68-0.68 $\mathbf{0.33}$ 0.90-0.90 $\mathbf{0.03}$  + $\mathbf{0.04}$ 0.43-0.43
+ NVSpeech 0.38-0.38 - 0.60-0.60 0.43-0.43 - 0.90-0.90
+ SMIIP-NV 0.11-0.11 - - 0.25-0.25 - -
+ SynParaSpeech 0.46-0.46 0.62-0.62 0.30-0.30 - - -
+ MNV-17 1.35-1.35 0.10-0.10 0.10-0.10 0.13-0.13 0.96-0.96 1.22-1.22
+ NVS(TBO) $\mathbf{0.05}$ 0.17-0.17 0.25-0.25 0.75-0.75 0.38-0.38 0.07-0.07
Chinese (ZH)
CosyVoice2 0.77-0.77 - 1.62-1.62 1.58-1.58 - 1.78-1.78
F5-TTS - - - - - -
+ CapSpeech 1.50-1.50 - - 2.20-2.20 - 1.80-1.80
+ NVTTS 0.50-0.50 0.68-0.68 0.57-0.57 0.03-0.03 0.34-0.34 0.05-0.05  +
+ NVSpeech 0.47-0.47 - 0.10-0.10 0.58-0.58 - 0.02-0.02
+ SMIIP-NV 0.18-0.18 - - 0.06-0.06 - -
+ SynParaSpeech 0.13-0.13 0.60-0.60 0.67-0.67 - - -
+ MNV-17 0.83-0.83 0.26-0.26 0.60-0.60 0.53-0.53 0.32-0.32 0.25-0.25
+ NVS(TBO) 0.03-0.03 0.20-0.20 0.62-0.62 0.05-0.05 0.40-0.40 0.25-0.25
  • Negative values indicate that F5-TTS fine-tuned with NVS (TSA) outperforms the corresponding system. Due to its high CER in Chinese, Dia is excluded from the IMOS ZH evaluation. [throat] denotes [throatclearing], and “-” indicates unsupported NV types for the current comparison system.
Table 10: Example alignment between reference (REF) and model hypothesis (HYP) with non-verbal tag [coughing].
@@ A.3 Experimental Setup for Non-Verbal Speech Understaning
  • Insertions (e.g., repeated what, after) and substitutions (e.g., he vs. they) are observed.
Appendix B NonVerbalSpeech-38K Pipeline
B.1 Details on Incorporating Non-Verbal Tags into Speech Content
@@
To support non-verbal speech generation and understanding, it is necessary to integrate the textual content with non-verbal tags. The related algorithm is presented in Algorithm 2, which implements the timestamp-based ordering (TBO) method. It is worth noting that state-of-the-art ASR models—including Whisper-Large-V3 (Radford et al. 2023) and Paraformer-zh (Gao et al. 2022)—offer word-level timestamps. In practice, however, we have observed that these timestamps exhibit varying degrees of misalignment, with Paraformer-zh showing particularly noticeable discrepancies. We try using forced alignment tools, such as Montreal Forced Aligner (McAuliffe et al. 2017), to improve alignment. However, due to insertion, substitution, and deletion errors in the transcriptions, no significant improvement is achieved; in some cases, the results are even worse.
Algorithm 2 Insert Non-Verbal Tag into Transcription
0: Non-verbal segment $n=[n_{s},n_{e}]$; non-verbal label $l$; timestamped word list $W=\{(s_{i},e_{i},w_{i})\}_{i=1}^{N}$
0: Updated word list with inserted non-verbal tags
[…]
5:if $i_{e}-i_{s}>1$ then
6:  # Spans multiple words
7:  return $W[0:i_{s}]+[s\_tag]+W[i_{s}:i_{e}]+[e\_tag]+W[i_{e}+1:N]$
8:else
9:  # Single word
[…]
11:end if
@@ Appendix C NonVerbalSpeech-38K Dataset
Table 11: Source distribution in the original data and the NVS-38K dataset
Source Radio Dramas Comedy Sketches Cartoon Variety Shows Short Plays
Crawled (hrs) 17,400 7,200 3,600 1,400 1,200
NVS-38K (#) 21,668 7,273 5,019 1,217 809
Source Speeches Documentaries Movies Audiobooks Toy Unboxing
Crawled (hrs) 1,079 600 500 263 9
NVS-38K (#) 158 105 1,090 1,375 4
  • “Crawled (hrs)” shows the duration of the original crawled data; “NVS-38K (#)” shows the final sample counts in the proposed NonVerbalSpeech-38K dataset.
Using the proposed NonVerbalSpeech-38K pipeline, we construct the NonVerbalSpeech-38K dataset from a large colillection of speech data crawled from diverse online sources (Table 11). This diversity ensures coverage of a wide range of natural speaking styles and, importantly, includes rich and varied non-verbal phonemes. Radio dramas, comedy sketches, and cartoons contribute the most data, while movies and audiobooks yield the highest processing conversion rates, suggesting richness in non-verbal content. The resulting durations mostly fall within the range of 3 to 30 seconds (Figure 5), primarily because the Emilia (He et al. 2024) pipeline segments speech clips into chunks of 3–30 seconds. This segmentation operation also explains the peak near 30 seconds. Some segments slightly exceeding 30 seconds arise during Step 3 of the NonVerbalSpeech-38K pipeline (Algorithm 1), where non-verbal segments that do not fall within any speech region but are within a 1-second distance threshold are merged with the nearest speech region to maintain contextual continuity.
Limitations: We observe a language imbalance in the current dataset, because the original source datasets are predominantly in Chinese, rather than due to limitations of our pipeline. In fact, our frame-level detection model is trained exclusively on English data, yet it demonstrates strong generalization performance on the Chinese test set. Moreover, since the majority of samples in our dataset are in Chinese, this further substantiates the robustness and cross-lingual generalization ability of the proposed detection model. Thanks to our fully automatic and scalable pipeline, further expansion of the dataset is feasible, and future work will continue to increase both its size and language coverage, helping to mitigate the current language imbalance.